# Learning Spatial-Semantic Features for Robust Video Object Segmentation

### v1

@@ 3.3 Discriminative Query Propagation
We note that updating the target query memory directly with entire object patches generated based on online predicted masks is ineffective as the predicted masks often cover background noise, reducing target distinctiveness and leading to accumulating errors over time. To propagate target queries effectively across frames, we update the target queries with the most distinctive feature of the target object.
@@
where $\mathcal{M}$ represents a binary mask of the target, and $\mathcal{F}[i]$ denotes the $i$-th channel activation of the feature map $\mathcal{F}$. We use the strategy in Cutie [14] to generate the target mask $\mathcal{M}$, which uses the predicted segmentation results in the previous frame.
Discriminative query update. Based on the discriminative target feature $\mathcal{Q}_{s}$ generated from a new target sample, we can update target queries by dynamically calculating the relationship between the salient query and salient pixel features in an additive manner, which can be formulated as:
$\mathcal{Q}_{\text{out }}=(\alpha\odot\frac{\mathcal{A}}{\|\mathcal{A}\|_{2}}+\mathcal{Q}_{s})\mathcal{W}_{\text{out }}+\mathcal{Q},$ (5)
where $\alpha$ is a learnable parameter to scale the matrix, and $\mathcal{A}$ denotes the correspondence between salient query and salient pixel features, which is computed as.
@@
The proposed discriminative query propagation scheme adaptively refines target queries with the most representative features, which helps deal with the challenges of dramatic appearance variations in long-term videos.
@@ 4.1 Experimental Settings
Pretrained Backbone & Baseline. We conduct different pre-trained weights obtained using various pre-training methods, including MAE [20], DINOV2 [34], and Depth Anything [48]. To validate the effectiveness of our proposed discriminative query propagation, we first choose ResNet50 as the backbone for a baseline model. The multi-scale features from the backbone are utilized in the Decoder for fine prediction. Soft aggregation operation is used to merge the predicted masks in the multi-target scenario.
Training. Our training settings are similar to Cutie [14]. Note that, apart from ResNet pre-training, we do not pre-train other models on static images and directly trained them on video datasets (YoutubeVOS [47] and DAVIS [36]). To enhance the performance of our model, we utilize the MEGA dataset constructed by Cutie, which includes the YouTubeVOS [47], DAVIS [36], OVIS [37], MOSE [17], and BURST [1] datasets. We sample eight frames to train the model, and three are randomly selected to train the matching process. For each sequence, we randomly choose at most three targets for training. Point supervision in loss is adopted to reduce the memory requirements.
For optimization, AdamW [25] optimizer is performed with a learning rate of 5e-5 and a weight decay of 0.5. We train the model for 125K with batch size 16 on the video dataset, and 195k on the MEGA dataset. For specific parameter settings, please refer to our supplementary materials. All our models are trained on 8 x NVIDIA V100 GPUs and tested on an NVIDIA V100 GPU.
Inference. Our feature and query memory are updated every 5th frame during the testing phase. For longer sequences, we employ a long-term fusion strategy [10] for updating. To enhance storage quality, we skip frames without targets and do not store them. The test input size contains two scales: 480 for fair comparison and 600/720 for further improvement.
Table 1: Ablation study. All the variants are trained on the YoutubeVOS and DAVIS datasets.
@@
XMem [10] (Baseline) 53.3 62.0 57.6 81.0 77.4 84.5 50.0 45.5 54.4 49.5 45.2 53.7 85.5 84.3 88.6 80.3 88.6
Cutie [14] 64.0 60.0 67.9 84.2 80.6 87.7 56.2 51.8 60.5 60.1 55.9 64.2 86.1 85.8 90.5 80.0 88.0
@@ 4.2 Ablation Study
Effect of different components. Table 1 provides a comprehensive analysis of the components of our method. The first two rows describe current methods upon which we build improvements. Cutie enhances performance by embedding a query transformer within the model.
+ Discriminative Query, which performs the discriminative query generation by adaptively refining target queries with the most representative features in the query transformer. The performance gains compared to Cutie demonstrate the advantages of target modeling of discriminative query for VOS, especially in long-term benchmarks.
+ ViT, which directly utilizing ViTDet [29] to incorporate pre-trained model features at multiple scales. The performance denotes that directly employing ViT to generate multi-scale features is minimally beneficial for VOS tasks.
+ Spatial, which applies spatial dependencies modeling in learning multi-scale detail features. It is clear that the proposed spatial dependencies modeling significantly boosts model performance in terms of $\mathcal{J}\&\mathcal{F}$ on all the VOS benchmarks. It validates the superiority of using spatial dependencies learning to conduct feature generation.
+ Semantic, which inserts semantic information to embed the multi-scale features. The performance in three benchmarks, MOSE, DAVIS 2017, and YouTubeVOS 2019, is further improved. These improvements validate the benefits of semantic embedding, which improves the semantic understanding of targets.
Effect of different pre-trained weights. Table 2 also compares the effects of different training parameters on segmentation performance. We conduct experiments using three different pre-training parameters, and the results indicate that better pre-training parameters lead to more significant improvements in the model, especially those from Depth Anything [48] (8-query in the table). The Depth Anything model implicitly contains depth information about the scene, which contributes to its stronger representation capability, consistent with the findings of the Depth Anything paper.
Effect of different query numbers. Table 2 shows the impact of different query numbers. Performance decreases as the number of queries increases. This is because more queries introduce additional background noise, given the fewer targets present in most VOS benchmarks, except MOSE [17].
Effect of adapters. Table 2 presents a performance comparison between different adapters and our method. We replace the fusion module with various adapters and fully fine-tune them on the MEGA dataset, as indicated in related research that full fine-tuning achieves optimal performance [8, 45]. The experimental results demonstrate that our method is efficient in training and outperforms the other two methods in effectiveness.
@@
ResNet [21] (Baseline) 64.2 60.3 68.1 85.2 81.8 88.5 57.4 53.3 61.5 86.3 86.2 91.0 79.9 88.3 86.5 86.2 90.7 80.6 88.8
MAE [20] 67.0 62.6 71.3 85.3 81.3 89.3 64.4 60.2 68.7 86.4 85.6 90.2 80.6 89.0 86.4 85.4 89.7 81.2 89.3
DINOv2 [34] 67.5 63.5 71.6 86.1 82.7 89.6 63.2 58.7 67.6 87.3 86.6 91.7 80.9 89.9 86.7 86.2 91.0 80.6 89.1
ViT-small [48] 64.3 60.0 68.5 85.3 81.5 89.1 55.9 51.6 60.1 86.0 85.3 90.4 79.9 88.5 86.3 85.3 90.1 80.6 89.1
@@
Ours (Adapter [8] ) 72.2 67.8 76.4 87.2 84.1 90.4 71.1 66.7 75.6 87.1 86.1 91.0 81.6 89.8 87.1 85.9 90.9 81.4 90.1
Ours (CoMer [45]) 73.2 69.1 77.3 87.0 83.2 90.8 73.4 68.8 77.9 87.3 87.1 92.2 80.7 89.4 87.4 87.0 91.7 81.3 89.6
@@
4.3 Evaluations with State-Of-The-Art Methods
DAVIS 2017 [36] is a benchmark that offers densely annotated, high-quality, full-resolution videos with multiple objects of interest. Table 3 shows that our method achieves favorable performance (86.7%) against state-of-the-art only trained on YoutubeVOS and DAVIS. The gains are primarily due to the injection of semantic information, which helps handle sequences with object occlusion. With more training data and improved inference size, our method performs best on the DAVIS test set (89.1%).
Table 3: Quantitative comparisons on the MOSE, LVOS, DAVIS 2017, YouTube-VOS 2018 & 2019 dataset. The best two results are shown in red and blue color. In the table, * denotes the models are pre-trained using the additional datasets, including static image datasets [38, 28, 52, 13, 43] and the large BL30K dataset, and + means using a larger input image size while testing.
@@ 4.3 Evaluations with State-Of-The-Art Methods
Trained on the YouTube VOS, and DAVIS datasets
STM [33] * - - - - - - - - - 79.4 79.7 84.2 72.8 80.9 - - - - - -
AFB-URR [30] * - - - - - - - - - 79.6 78.8 83.1 74.1 82.6 - - - - - -
CFBI [49] * - - - 50.0 45.0 55.1 74.8 71.1 78.5 81.4 81.1 85.8 75.3 83.4 81.0 80.6 85.1 75.2 83.0 2.9
MiVOS [12]* - - - - - - 78.6 74.9 82.2 82.6 81.1 85.6 77.7 86.2 82.4 80.6 84.7 78.1 86.4 -
STCN [11] * 52.5 48.5 56.6 45.8 41.1 50.5 77.8 74.3 81.3 84.3 83.2 87.9 79.0 87.3 84.2 82.6 87.0 79.4 87.7 13.2
Swin-B-AOT-L [50] * 59.4 53.6 65.2 59.4 53.6 65.2 81.2 77.3 85.1 85.1 85.1 90.1 78.4 86.9 85.3 84.6 89.5 79.3 87.7 12.1
DeAOT-R50 [51] 59.0 54.6 63.4 - - - 80.7 76.9 84.5 86.0 84.9 89.9 80.4 88.7 85.6 84.2 89.2 80.2 88.8 11.7
XMem - - - - - - 79.8 76.3 83.4 84.3 83.9 88.8 77.7 86.7 84.2 83.8 88.3 78.1 86.7
XMem[10] * 53.3 62.0 57.6 50.0 45.5 54.4 81.0 77.4 84.5 85.7 84.6 89.3 80.2 88.7 85.5 84.3 88.6 80.3 88.6 22.6
ISVOS [42] * - - - - - - 82.8 79.3 86.2 86.3 85.5 90.2 80.5 88.8 86.1 85.2 89.7 80.7 88.9 5.8
SimVOS-B [44] 61.6 57.9 65.3 - - - 82.3 78.7 85.8 - - - - - 84.2 83.1 87.5 79.1 87.2 3.3
Cutie [14]* 64.0 60.0 67.9 - - - 84.2 80.6 87.7 86.1 85.5 90.0 80.6 88.3 86.1 85.8 90.5 80.0 88.0 36.4
Cutie [14]*+ 66.2 62.3 70.1 56.2 51.8 60.5 85.9 82.6 89.2 - - - - - 86.9 86.2 90.7 81.6 89.2 17.9
JointFormer [32] - - - - - - 87.0 83.4 90.6 86.0 86.0 91.0 79.5 87.5 86.2 85.7 90.5 80.4 88.2
JointFormer [32]* - - - - - - 88.1 84.7 91.6 87.6 86.4 91.0 82.2 90.7 87.4 86.5 90.9 82.0 90.3 3.0
Ours 68.5 64.5 72.6 - - - 86.7 82.7 90.8 87.4 87.0 92.0 80.9 89.7 87.5 86.8 91.8 81.3 89.9 8.0
Ours+ 70.5 66.5 74.6 66.5 62.1 70.8 87.8 84.6 91.3 87.6 86.9 91.7 81.5 90.1 87.7 86.8 91.6 82.2 90.8 4.5
@@
DEVA [15] 66.5 62.3 70.8 54.0 49.0 59.0 83.2 79.6 86.8 86.2 85.4 89.9 80.5 89.1 85.8 84.8 89.2 80.3 88.8 25.3
Cutie [14] * 69.9 65.8 74.1 - - - 86.1 82.4 89.9 87.0 86.4 91.1 81.4 89.2 87.0 86.0 90.5 82.0 89.6 36.4
Cutie [14]*+ 71.7 67.6 75.8 66.7 62.4 71.0 88.1 84.7 91.4 - - - - - 87.5 86.3 90.6 82.7 90.5 17.9
Ours 74.0 69.8 78.3 - - - 87.8 84.0 91.7 88.0 87.0 91.8 82.5 90.7 88.1 87.4 92.5 81.9 90.7 8.0
Ours+ 75.1 71.0 79.2 73.0 68.3 77.8 89.1 85.8 92.4 88.5 87.6 92.6 82.7 91.3 88.5 87.3 92.0 83.1 91.4 4.5
YouTubeVOS 2018 [47] has 3471 videos with 65 categories for training and videos for validation. In the validation, there are 26 categories that the model has not seen in its training, enabling us to evaluate its generalization ability for 350 class-agnostic targets. Table 3 illustrates that our method achieves better performance on seen and unseen categories than previous state-of-the-art models. The performance gain is attributed to spatial-semantic feature representation and representative target correlation.
YouTubeVOS 2019 is an extension of YouTube-VOS 2018, featuring a greater number of masked targets and including more challenging sequences in its val set. In Table 3, our approach exhibits competitive performance relative to the leading state-of-the-art methods. Remarkably, our model achieves the highest performance (87.5%) without relying on pre-training. This underscores the effectiveness of our method in leveraging inherent semantic-spatial information and discriminative query representation.
MOSE [17] contains 2149 video clips and 5200 objects from 36 categories under more complex scenes. Compared to existing state-of-the-art methods, our model gains great improvement in this dataset. The benefits primarily stem from the discriminative query capturing more representative target information. Additionally, the feature extraction module provides matching with more semantic and spatially informative features.
LVOS [22] is a new long-term dataset to validate the robustness of VOS methods when facing these challenges. It contains videos 20 times longer than those in the existing VOS datasets, typically featuring a single target per video. In this benchmark, our method achieves a performance gain (66.5% compared to 56.2%). This improvement is attributed to the superior quality of memorized features and queries, which enhance the robustness of our approach in long-term video scenarios. Our method demonstrates enhanced stability and effectiveness in long-duration contexts by incorporating more discriminative queries that capture crucial target information into the object memory.
Results with larger test size. Following Cutie, we notice that the performance of our method can be further improved by changing hyperparameters, such as test input scales, memory interval, memorize or not, and the size of the memory bank during inference. As is shown in Table 3, owing to the robust spatial-semantic feature learning and discriminative query propagation, our method achieves state-of-the-art performance on most benchmarks (e.g. 75.1% on MOSE [17] val set). Besides, our method is only trained with video data, without any pre-train, but still achieves better performance.
@@
Visualized Results. We visualize some challenging sequences, including scenarios involving part-to-whole changing and faint objects, as well as those where targets closely resemble the background. In Figure 3(a), our approach shows favorable segmentation performance against the state-of-the-art methods. For instance, our model accurately segments targets and provides detailed results in a video that includes gorillas, necks, and a riding man. This is attributed to our proposed discriminative query propagation, which enables our model to capture more detailed target information by focusing on the most representative features, thereby facilitating precise segmentation in these challenging scenarios Additionally, incorporating semantic features in our model ensures the model understands the semantic prior of the interested targets, such as in the case of the sequence of man riding from YoutubeVOS 2019.
@@
Figure 3(b) shows the segmentation and tracking performance on long-term samples with fast motion and dramatic appearance variations. In this sequence, two similar dogs continuously change their position and posture, which is challenging for the model to identify and track them accurately. Our VOS method performs accurately in tracking and segmenting the targets, whereas Cutie fails to identify and segment the targets in this scenario. Good performance in scenarios like this is due to the semantic-aware feature and discriminative query representation learned by our model. More visualization and videos can be found in the appendix.
@@
Figure 3: Visualized results on challenging scenarios. The proposed method performs well on these challenging sequences, while other methods suffer from semantic understanding and thus cannot segment the whole objects. (b) Our method handles the challenging sequence with fast-motion objects better than Cutie [14] .
5 Limitations

### v2

@@ 3.3 Discriminative Query Propagation
To propagate target queries effectively across frames, we update the target queries with the most distinctive feature of the target object.
@@
where $\odot$ denotes the dot product, $\mathcal{M}$ represents a binary mask of the target, and $\mathcal{F}[i]$ denotes the $i$-th channel activation of the feature map $\mathcal{F}$. We use the strategy in Cutie (Cheng et al., 2023a) to generate the target mask $\mathcal{M}$, which uses the predicted segmentation results in the previous frame.
Query propagation. Based on the discriminative target feature $\mathcal{Q}_{s}$ generated from the target sample and the input target query $\mathcal{Q}_{\text{in}}$ at the current stage, the output query $\mathcal{Q}_{\text{out}}$ is propagated as:
$\mathcal{Q}_{\text{out}}=(\alpha\odot\frac{\mathcal{A}}{\|\mathcal{A}\|_{2}}+\mathcal{Q}_{s})\mathcal{W}_{\text{out }}+\mathcal{Q_{\text{in}}},$ (5)
where $\alpha$ is a learnable parameter to scale the matrix, $\mathcal{W}_{\text{out }}$ is a learnable projection matrix, and $\mathcal{A}$ denotes the correspondence between salient query and salient pixel features, which is computed as.
@@
The proposed discriminative query propagation scheme adaptively refines target queries with the most representative features, which helps deal with the challenges of dramatic appearance variations in long-term videos. We also present the similarity map between the discriminative query and the feature in the last two rows, which indicates that our proposed discriminative query can represent and distinguish different objects.
@@ 4.1 Experimental Settings
Network. The full version of the proposed method uses a pre-trained ViT-base (Dosovitskiy et al., 2021) model as the ViT branch and constructs a simple CNN module for multi-scale feature generation. The Spatial-Semantic block performs feature modeling after every block of the ViT branch. Upon the target association part, the Decoder predicts target masks based on multi-scale features. We use a soft aggregation operation for the scenario containing multiple target objects to merge the predicted masks.
In all experiments, we set the number of Spatial-Semantic Blocks to N=4.
Training. We adopt similar training settings as those of Cutie (Cheng et al., 2023a). Note that, apart from ResNet pre-training, we do not pre-train other models on static images and directly trained them on video datasets (YoutubeVOS (Xu et al., 2018) and DAVIS (Pont-Tuset et al., 2017)). To enhance the performance of our model, we use the MEGA dataset constructed by Cutie, which includes the YouTubeVOS (Xu et al., 2018), DAVIS (Pont-Tuset et al., 2017), OVIS (Qi et al., 2022), MOSE (Ding et al., 2023) and BURST (Athar et al., 2023) datasets. We sample $8$ frames to train the model and randomly select 3 of them to train the matching process. For each sequence, we randomly choose at most 3 targets for training. Point supervision in loss is adopted to reduce the memory requirements.
For optimization, the AdamW (Kingma & Ba, 2014) optimizer is used with a learning rate of 5e-5 and a weight decay of 0.5. We train the model for 125K with a batch size of 16 on the video dataset, and 195k on the MEGA dataset. For specific parameter settings, please refer to our supplementary material. All our models are trained on a machine with 8 x NVIDIA V100 GPUs and tested using one NVIDIA V100 GPU.
Inference. Our feature and query memories are updated every 5th frame during the testing phase. For longer sequences, we employ a long-term fusion strategy (Cheng & Schwing, 2022) for updating. We skip frames without targets. We use two kinds of input sizes, 480 and 600/720, for more detailed comparisons. In this section, we compare all methods with input size 480p.
Table 1: Ablation study. All the variants are well aligned using the same settings and trained on the YoutubeVOS and DAVIS datasets. S3 denotes the proposed algorithm.
@@
XMem (Cheng & Schwing, 2022) (Baseline) 53.3 62.0 57.6 81.0 77.4 84.5 50.0 45.5 54.4 49.5 45.2 53.7 85.5 84.3 88.6 80.3 88.6
Cutie (Cheng et al., 2023a) 64.0 60.0 67.9 84.2 80.6 87.7 56.2 51.8 60.5 60.1 55.9 64.2 86.1 85.8 90.5 80.0 88.0
@@ 4.2 Ablation Study
Effect of every component. Table 1 provides a comprehensive analysis of the components of our method. The first two rows describe the current methods upon which we build improvements. Cutie improves performance by using a query transformer within the model.
Baseline. We adopt XMem as the baseline model, a representative VOS framework with a well-designed memory mechanism.
+ Discriminative Query, which performs discriminative query generation by adaptively refining target queries with the most representative features of the query transformer. To validate the effectiveness of our proposed discriminative query propagation, we first choose ResNet50 as the backbone for feature generation (simply comparing query propagation with Cutie (Cheng et al., 2023a)). Performance gains compared to Cutie demonstrate the advantages of target modeling of discriminative queries for VOS, especially in long-term benchmarks.
+ ViT, which directly utilizes ViTDet (Li et al., 2022b) to incorporate pre-trained model features at multiple scales. The performance indicates that directly employing ViT to generate multi-scale features is minimally beneficial for VOS tasks.
+ Spatial-semantic, which applies the spatial-semantic blocks to integrate semantic information and multi-scale features. It is clear that the proposed spatial dependency modeling significantly boosts model performance in terms of $\mathcal{J}\&\mathcal{F}$ on all VOS benchmarks. These improvements validate the benefits of the spatial-semantic network, which improves the semantic understanding of targets.
Effect of different pre-trained weights. Table 2 also compares the performance of using different training parameters. We conduct experiments using three different pre-training parameters, and the results indicate that better pre-training parameters lead to more significant improvements in the model, especially those from Depth Anything (Yang et al., 2024) (8-query in Table 2). The Depth Anything model implicitly contains depth information about the scene, which contributes to its stronger representation capability, consistent with the findings of the Depth Anything paper.
Effect of different query numbers. Table 2 shows that performance drops as the number of queries increases. This is because more queries introduce additional background noise, given the fewer targets in most VOS benchmarks, except MOSE (Ding et al., 2023).
Effect of adapters. Table 2 presents a performance comparison between different adapters and our method. We replace the fusion module with various adapters and fully fine-tune them on the MEGA dataset, as indicated in related research that full fine-tuning achieves optimal performance (Chen et al., 2022; Xia et al., 2024). The experimental results demonstrate that our method is efficient in training and outperforms the other two methods in accuracy.
@@
ResNet (He et al., 2016) (Baseline) 64.2 60.3 68.1 85.2 81.8 88.5 57.4 53.3 61.5 86.3 86.2 91.0 79.9 88.3 86.5 86.2 90.7 80.6 88.8
MAE (He et al., 2022) 67.0 62.6 71.3 85.3 81.3 89.3 64.4 60.2 68.7 86.4 85.6 90.2 80.6 89.0 86.4 85.4 89.7 81.2 89.3
DINOv2 (Oquab et al., 2023) 67.5 63.5 71.6 86.1 82.7 89.6 63.2 58.7 67.6 87.3 86.6 91.7 80.9 89.9 86.7 86.2 91.0 80.6 89.1
ViT-small (Yang et al., 2024) 64.3 60.0 68.5 85.3 81.5 89.1 55.9 51.6 60.1 86.0 85.3 90.4 79.9 88.5 86.3 85.3 90.1 80.6 89.1
Depth_Anything 68.5 64.5 72.6 86.7 82.7 90.8 66.5 62.1 70.8 87.4 87.0 92.0 80.9 89.7 87.5 86.8 91.8 81.3 89.9
@@
Ours (Adapter (Chen et al., 2022) ) 72.2 67.8 76.4 87.2 84.1 90.4 71.1 66.7 75.6 87.1 86.1 91.0 81.6 89.8 87.1 85.9 90.9 81.4 90.1
Ours (CoMer (Xia et al., 2024)) 73.2 69.1 77.3 87.0 83.2 90.8 73.4 68.8 77.9 87.3 87.1 92.2 80.7 89.4 87.4 87.0 91.7 81.3 89.6
@@
4.3 Comparisons with State-of-the-Art Methods
DAVIS 2017 (Pont-Tuset et al., 2017) is a benchmark that offers densely annotated, high-quality, full-resolution videos with multiple objects of interest. Table 3 shows that our method achieves favorable performance (86.7%) against state-of-the-art methods when trained on YoutubeVOS and DAVIS, without any pre-training. The gains are primarily due to the injection of semantic information, which helps handle sequences with object occlusion. When using the MEGA dataset as the training data, our method performs best on the DAVIS test set (87.8%).
YouTubeVOS 2018 (Xu et al., 2018) contains 347 videos of 65 categories for training and videos for validation. For validation, there are 26 categories that the model has not seen in its training, allowing us to evaluate its generalization ability for 350 class-agnostic targets. Table 3 illustrates that our method performs better on seen and unseen categories than existing state-of-the-art models. The performance gain can be attributed to spatial-semantic feature representation and representative target correlation.
YouTubeVOS 2019 is an extension of YouTube-VOS 2018, featuring a larger number of masked targets and more challenging sequences in its validation set. In Table 3, our approach exhibits competitive performance compared to the leading state-of-the-art methods. Remarkably, our model achieves the highest performance (87.5%) without relying on pre-training. This underscores the effectiveness of our method in leveraging inherent semantic-spatial information and discriminative query representation.
MOSE (Ding et al., 2023) contains 2149 video clips and 5200 objects from 36 categories under more complex scenes. Compared to existing state-of-the-art methods, our model gains great improvement in this dataset. The benefits stem primarily from the discriminative query that captures more representative target information. Additionally, the feature extraction module provides matching with more semantic and spatially informative features.
LVOS (Hong et al., 2023) is a new long-term dataset to validate the robustness of VOS methods when faced with these challenges. It contains videos 20 times longer than existing VOS datasets, typically featuring a single target per video. In this benchmark, our method achieves a performance gain (66.5% compared to 56.2%). This improvement is attributed to the quality of memorized features and queries, which enhances the robustness of our approach in long-term video scenarios. Our method demonstrates enhanced stability and effectiveness in long-duration contexts by incorporating more discriminative queries that capture crucial target information into the object memory.
Comparison with Segment Anything 2 (Ravi et al., 2024). Table 4 shows the detailed comparison results between our method and SAM2. Specifically, SAM2 used a large-scale mixed training dataset (comprising 1. 3% DAVIS, approximately 9.4% MOSE, approximately 9.2% YouTubeVOS, approximately 15.5% SA-1B, approximately 49.5% SA-V, and approximately 15. 1% Internal, approximately 120.6k videos). In contrast, our model is trained solely on VOS datasets (DAVIS, MOSE, YouTubeVOS, OVIS, and BURST, about 11.4K videos). To ensure a relatively fair comparison, we use the same model size (ViTb) and the same input image size (1024p). In all benchmarks evaluated, our method outperforms SAM2 in all metrics, especially on the challenging LVOSV2 val dataset (+7.9%). Even when SAM2 (Ravi et al., 2024) uses Hiera-L as a feature extractor, our model demonstrates comparable results, especially on LVOS V2 (Hong et al., 2023) val dataset (83.7% vs 78.1%).
Table 3: Quantitative comparisons on the MOSE, LVOS, DAVIS 2017, YouTube-VOS 2018 & 2019 dataset. The best two results are shown in red and blue color. In the table, * denotes that the models are pre-trained using the static image datasets (Shi et al., 2016; Li et al., 2020; Zeng et al., 2019; Cheng et al., 2020; Wang et al., 2017). In the Appendix, we compare methods with external training data (e.g., large BL30K dataset for pre-training) and improved test size (600p/720p).
@@ 4.3 Comparisons with State-of-the-Art Methods
Trained only on the YouTube VOS, and DAVIS datasets
MiVOS (Cheng et al., 2021b)* - - - - - - 78.6 74.9 82.2 82.6 81.1 85.6 77.7 86.2 82.4 80.6 84.7 78.1 86.4 -
STCN (Cheng et al., 2021a) * 52.5 48.5 56.6 45.8 41.1 50.5 77.8 74.3 81.3 84.3 83.2 87.9 79.0 87.3 84.2 82.6 87.0 79.4 87.7 13.2
Swin-B-AOT-L (Yang et al., 2021) * 59.4 53.6 65.2 54.4 49.3 59.4 81.2 77.3 85.1 85.1 85.1 90.1 78.4 86.9 85.3 84.6 89.5 79.3 87.7 12.1
DeAOT-R50 (Yang & Yang, 2022) 59.0 54.6 63.4 - - - 80.7 76.9 84.5 86.0 84.9 89.9 80.4 88.7 85.6 84.2 89.2 80.2 88.8 11.7
XMem (Cheng & Schwing, 2022) - - - - - - 79.8 76.3 83.4 84.3 83.9 88.8 77.7 86.7 84.2 83.8 88.3 78.1 86.7
XMem (Cheng & Schwing, 2022) * 53.3 62.0 57.6 50.0 45.5 54.4 81.0 77.4 84.5 85.7 84.6 89.3 80.2 88.7 85.5 84.3 88.6 80.3 88.6 22.6
ISVOS (Wang et al., 2023) * - - - - - - 82.8 79.3 86.2 86.3 85.5 90.2 80.5 88.8 86.1 85.2 89.7 80.7 88.9 5.8
SimVOS-B (Wu et al., 2023) 61.6 57.9 65.3 - - - 82.3 78.7 85.8 - - - - - 84.2 83.1 87.5 79.1 87.2 3.3
Cutie (Cheng et al., 2023a)* 64.0 60.0 67.9 56.2 51.8 60.5 84.2 80.6 87.7 86.1 85.5 90.0 80.6 88.3 86.1 85.8 90.5 80.0 88.0 36.4
JointFormer (Zhang et al., 2023) - - - - - - 87.0 83.4 90.6 86.0 86.0 91.0 79.5 87.5 86.2 85.7 90.5 80.4 88.2 3.0
JointFormer (Zhang et al., 2023)* - - - - - - 87.6 84.2 91.1 87.0 86.2 91.0 81.4 89.3 87.0 86.1 90.6 82.0 89.5 3.0
S3 (Ours) 68.5 64.5 72.6 66.5 62.1 70.8 86.7 82.7 90.8 87.4 87.0 92.0 80.9 89.7 87.5 86.8 91.8 81.3 89.9 13.1
@@
DEVA (Cheng et al., 2023b) 66.5 62.3 70.8 54.0 49.0 59.0 83.2 79.6 86.8 86.2 85.4 89.9 80.5 89.1 85.8 84.8 89.2 80.3 88.8 25.3
Cutie (Cheng et al., 2023a) * 69.9 65.8 74.1 66.7 62.4 71.0 86.1 82.4 89.9 87.0 86.4 91.1 81.4 89.2 87.0 86.0 90.5 82.0 89.6 36.4
S3 (Ours) 74.0 69.8 78.3 73.0 68.3 77.8 87.8 84.0 91.7 88.0 87.0 91.8 82.5 90.7 88.1 87.4 92.5 81.9 90.7 13.1
Table 4: Comparasion with SAM 2 (Ravi et al., 2024). It shows that our method performs favorably against SAM 2.
Dataset LVOSV2-val LVOS-val DAVIS17-test MOSE-val YouTube-VOS 2019 val
[…]
SAM 2 (Hiera-B+) 75.8 72.0 79.6 74.9 70.2 79.6 88.3 85.0 91.5 75.8 71.8 79.9 88.4 85.2 91.6
SAM 2 (Hiera-L) 78.1 74.3 81.9 76.1 71.6 80.6 89.0 85.8 92.2 77.2 73.3 81.2 89.1 86.0 92.2
S3 (ViTb+) (+7.9) 83.7 80.1 87.2 (+1.1) 76.0 71.2 80.8 (+1.5) 89.8 86.4 93.1 (+0.1) 75.9 71.8 79.8 (+0.3) 88.7 85.5 91.9
@@
Visualized Results. We visualize some challenging sequences, including scenarios involving part-to-whole-changing and faint objects and those where targets closely resemble the background. In Figure 4(a), our approach shows favorable segmentation performance against the state-of-the-art methods. For example, our model accurately segments targets and provides detailed results in a video that includes gorillas, necks, and a riding man. This can be attributed to our proposed discriminative query propagation, which enables our model to capture more detailed target information by focusing on the most representative features, thereby facilitating precise segmentation in these challenging scenarios Additionally, incorporating semantic features in our model ensures the model understands the semantic prior of the interested targets, such as in the case of the sequence of man riding from YoutubeVOS 2019.
@@
Figure 4(b) shows the segmentation and tracking performance on long-term sequences with fast motion and dramatic appearance variations. In this sequence, two similar dogs continuously change their position and posture, making it challenging for the model to identify and track them accurately. Our VOS method performs accurately in tracking and segmenting targets, whereas Cutie fails to identify and segment the targets in this scenario. Good performance in scenarios like this is due to the semantic-aware feature and discriminative query representation learned by our model. More visualization and videos can be found in the Supplementary Material.
@@
Figure 4: Visualized results on challenging scenarios. The proposed method performs well on these challenging sequences, while other methods suffer from semantic understanding and thus cannot segment the whole objects. (b) Our method handles the challenging sequence with fast-motion objects better than Cutie (Cheng et al., 2023a) .
4.4 Limitations